The probability that a region prone to hurricanes will be hit by a hurricane in any single year is 15. What is the probability of a hurricane at least once in the next five years?
Answer by stanbon(75887)

The probability that a region prone to hurricanes will be hit by a hurricane in any single year is 0.15.
What is the probability of a hurricane at least once in the next five years
-----
P(at least one) = 1 - P(none in five years)
-----
= 1 - 0.85^5
----
= 1 - 0.44
---
= 0.56
================
Cheers,
Stan H.
================